OpenStudy (anonymous):

Evaluate the expression if a = 2.

OpenStudy (anonymous):



OpenStudy (anonymous):

\[5a ^{4}\]

OpenStudy (anonymous):

\[2^4=2*2*2*2=4*4=16\] times 5 \[16*5=80\]